# Magneto-optical transport in tilted type-I double-Weyl semimetals in the presence of orbital magnetic moment

### v1

@@ III.3 Calculatons of quadractic-B contributon to the conductivity $\sigma_{ab}^{(B^{2})}$
Substituting Eq.(III) with B terms up to second order into the first term of Eq.(III), we get conductivities components
@@
$\sigma_{ab}^{(B^{2},\Omega)}(\omega)=\frac{\tau}{\hbar^{2}(1-i\omega\tau)}\frac{e^{4}}{(2\pi)^{3}}\int d^{3}k[\mathit{v}_{a}^{s}\mathit{v}_{b}^{s}(\Omega_{\bm{k}}^{s}\cdot\textbf{B})^{2}-(\mathit{v}_{a}^{s}B_{b}+\mathit{v}_{b}^{s}B_{a})(\Omega_{\bm{k}}^{s}\cdot\bm{\mathit{v}}_{\bm{k}}^{s})(\Omega_{\bm{k}}^{s}\cdot\textbf{B})+B_{a}B_{b}(\Omega_{\bm{k}}^{s}\cdot\bm{\mathit{v}}_{\bm{k}}^{s})^{2}]\Bigl{(}-\frac{\partial f_{0}^{s}}{\partial\epsilon_{\bm{k}}^{s}}\Bigr{)}$
$\sigma_{ab}^{(B^{2},m)}(\omega)=\frac{\tau}{\hbar^{2}(1-i\omega\tau)}\frac{e^{3}}{(2\pi)^{3}}\int d^{3}k\Bigl{[}\frac{\partial}{\partial\textbf{k}}\cdot[\mathit{v}_{a}^{s}B_{b}\Omega_{\bm{k}}^{s}](\textbf{m}_{\bm{k}}^{s}\cdot\textbf{B})-\frac{\partial[\mathit{v}_{a}^{s}(\Omega_{\bm{k}}^{s}\cdot\textbf{B})]}{\partial k_{b}}(\textbf{m}_{\bm{k}}^{s}\cdot\textbf{B})+\frac{\partial[B_{a}^{s}(\Omega_{\bm{k}}^{s}\cdot\mathit{\bm{v}}_{\bm{k}}^{s})]}{\partial k_{b}}(\textbf{m}_{\bm{k}}^{s}\cdot\textbf{B})$
$+\frac{\partial(\textbf{m}_{\bm{k}}^{s}\cdot\textbf{B})}{\partial k_{a}}(\Omega_{\bm{k}}^{s}\cdot\textbf{B})\mathit{v}_{b}^{s}-\frac{\partial(\textbf{m}_{\bm{k}}^{s}\cdot\textbf{B})}{\partial k_{a}}(\Omega_{\bm{k}}^{s}\cdot\mathit{\bm{v}}_{\bm{k}}^{s})B_{b}^{s}-\frac{\partial^{2}(\textbf{m}_{\bm{k}}^{s}\cdot\textbf{B})}{e\partial k_{a}\partial k_{b}}(\textbf{m}_{\bm{k}}^{s}\cdot\textbf{B})-B_{a}\frac{\partial(\textbf{m}_{\bm{k}}^{s}\cdot\textbf{B})}{\partial\bm{k}}\cdot\Omega_{\bm{k}}^{s}\mathit{v}_{b}^{s}\Bigr{]}\Bigl{(}-\frac{\partial f_{0}^{s}}{\partial\epsilon_{\bm{k}}^{s}}\Bigr{)}$
$\sigma_{ab}^{(B^{2},\Omega)}(\omega)$ and $\sigma_{ab}^{(B^{2},m)}(\omega)$ Berry-curvature $\Omega_{\bm{k}}^{s}$ and the orbital magnetic moment $\textbf{m}_{\bm{k}}^{s}$ dependent conductivities respectively. Consider the fllowing two cases
@@
In this case, Hall conductivities components are zero and one can get following expressions for longtudinal components
$\sigma_{zz}^{(B^{2},\Omega)}(\omega)=\frac{5\pi\mathit{v}_{F}^{2}\hbar}{4\mu\alpha_{2}}\sigma_{0}^{(B^{2})}$
$\sigma_{xx}^{(B^{2},\Omega)}(\omega)=8\sigma_{1}^{(B^{2})}\Bigl{[}\frac{-2(15-25t_{s}^{2}+8t_{s}^{4})}{t_{s}^{6}}$
$-15\frac{(1-t_{s}^{2})^{2}}{t_{s}^{7}}\ln\frac{1-t_{s}}{1+t_{s}}\Bigr{]}$ (37)
@@
$\sigma_{xx}^{(B^{2},m)}(\omega)=-\frac{\pi\mathit{v}_{F}^{2}\hbar}{\mu\alpha_{2}}\sigma_{0}^{(B^{2})}$ (38)
$\sigma_{zz}^{(B^{2},m)}(\omega)=\frac{16}{5}\sigma_{1}^{(B^{2})}\Bigl{[}\frac{(30-20t_{s}^{2}+t_{s}^{4})}{t_{s}^{6}}$
$+\frac{5}{2}\frac{(6-t_{s}^{2}+t_{s}^{4})}{t_{s}^{7}}\ln\frac{1-t_{s}}{1+t_{s}}]$ (39)
where $\sigma_{1}^{(B^{2})}=\frac{e^{2}\tau(eB)^{2}\alpha^{2}}{8\pi^{2}\hbar^{2}(1-i\omega t)\hbar\mathit{v}_{F}}$.
Case-II In this case both longitudinal components as well as planer Hall conductivities are non-zero. The longtudnal components have the following expressions
@@
$\sigma_{xx}^{(B^{2},\Omega)}(\omega)$ $=$ $\frac{\pi\mathit{v}_{F}^{2}\hbar}{32\mu\alpha_{2}}\sigma_{1}^{(B^{2})}[(31+16t_{s}^{2})\cos^{2}\gamma+5\pi^{2}\sin^{2}\gamma]$
@@
$\sigma_{zz}^{(B^{2},\Omega)}(\omega)$ $=$ $\sigma_{1}^{(B^{2})}\frac{\mathit{v}_{F}^{4}\hbar^{2}}{15\mu^{2}}(1+7t_{s}^{2})$
@@
$\sigma_{xy}^{(B^{2},\Omega)}(\omega)=\sigma_{1}^{(B^{2})}\frac{\pi\mathit{v}_{F}^{2}\hbar}{16\mu\alpha_{2}}(13+8t_{s}^{2})\sin\gamma\cos\gamma$ (42)
The magnetic orbital moment contrbutions of conductivities
$\sigma_{xx}^{(B^{2},m)}(\omega)=\frac{\pi\mathit{v}_{F}^{2}}{2\mu\alpha_{2}}\sigma_{0}^{(B^{2})}(2\cos^{2}\gamma-\sin^{2}\gamma)$ (43)
$\sigma_{zz}^{(B^{2},m)}(\omega)=-\frac{\mathit{v}_{F}^{4}\hbar^{2}}{15\mu^{2}\alpha_{2}^{2}}\sigma_{1}^{(B^{2})}(1-t_{s}^{2})$ (44)
@@
$\sigma_{xy}^{(B^{2},m)}(\omega)=\frac{2\pi^{2}\mathit{v}_{F}^{2}\hbar}{\mu\alpha_{2}}\sin\gamma\cos\gamma\sigma_{1}^{(B^{2})}$ (45)

### v2

@@ III.3 Calculations of quadratic-B contribution to the conductivity $\sigma_{ab}^{(B^{2})}$
Substituting Eq.(16) with B terms up to second order into the first term of Eq.(17), we get conductivities components
@@
$\begin{split}\sigma_{ab}^{(B^{2},\Omega)}(\omega)=&\frac{\tau}{\hbar^{2}(1-i\omega\tau)}\frac{e^{4}}{(2\pi)^{3}}\int d^{3}k[\mathit{v}_{a}^{s}\mathit{v}_{b}^{s}(\Omega_{\bm{k}}^{s}\cdot\textbf{B})^{2}-(\mathit{v}_{a}^{s}B_{b}+\mathit{v}_{b}^{s}B_{a})(\Omega_{\bm{k}}^{s}\cdot\bm{\mathit{v}}_{\bm{k}}^{s})(\Omega_{\bm{k}}^{s}\cdot\textbf{B})\\ &+B_{a}B_{b}(\Omega_{\bm{k}}^{s}\cdot\bm{\mathit{v}}_{\bm{k}}^{s})^{2}]\Bigl{(}-\frac{\partial f_{0}^{s}}{\partial\epsilon_{\bm{k}}^{s}}\Bigr{)}\end{split}$ (52)
$\begin{split}\sigma_{ab}^{(B^{2},m)}(\omega)=&\frac{\tau}{\hbar^{2}(1-i\omega\tau)}\frac{e^{3}}{(2\pi)^{3}}\int d^{3}k\Bigl{[}\frac{\partial}{\partial\textbf{k}}\cdot[\mathit{v}_{a}^{s}B_{b}\Omega_{\bm{k}}^{s}](\textbf{m}_{\bm{k}}^{s}\cdot\textbf{B})-\frac{\partial[\mathit{v}_{a}^{s}(\Omega_{\bm{k}}^{s}\cdot\textbf{B})]}{\partial k_{b}}(\textbf{m}_{\bm{k}}^{s}\cdot\textbf{B})+\frac{\partial[B_{a}^{s}(\Omega_{\bm{k}}^{s}\cdot\mathit{\bm{v}}_{\bm{k}}^{s})]}{\partial k_{b}}(\textbf{m}_{\bm{k}}^{s}\cdot\textbf{B})+\\ &\frac{\partial(\textbf{m}_{\bm{k}}^{s}\cdot\textbf{B})}{\partial k_{a}}(\Omega_{\bm{k}}^{s}\cdot\textbf{B})\mathit{v}_{b}^{s}-\frac{\partial(\textbf{m}_{\bm{k}}^{s}\cdot\textbf{B})}{\partial k_{a}}(\Omega_{\bm{k}}^{s}\cdot\mathit{\bm{v}}_{\bm{k}}^{s})B_{b}^{s}-\frac{\partial^{2}(\textbf{m}_{\bm{k}}^{s}\cdot\textbf{B})}{e\partial k_{a}\partial k_{b}}(\textbf{m}_{\bm{k}}^{s}\cdot\textbf{B})-B_{a}\frac{\partial(\textbf{m}_{\bm{k}}^{s}\cdot\textbf{B})}{\partial\bm{k}}\cdot\Omega_{\bm{k}}^{s}\mathit{v}_{b}^{s}\Bigr{]}\Bigl{(}-\frac{\partial f_{0}^{s}}{\partial\epsilon_{\bm{k}}^{s}}\Bigr{)}\end{split}$ (53)
$\sigma_{ab}^{(B^{2},\Omega)}(\omega)$ and $\sigma_{ab}^{(B^{2},m)}(\omega)$ Berry-curvature $\Omega_{\bm{k}}^{s}$ and the orbital magnetic moment $\textbf{m}_{\bm{k}}^{s}$ dependent conductivities respectively. Consider the following two cases
@@
In this case, Hall conductivities components are zero and one can get the following expressions for longitudinal components
for J=1
$\sigma_{zz}^{(B^{2},\Omega)}(\omega)$ $=$ $\frac{\mathit{v}_{F}^{3}}{15\pi\mu^{2}}\sigma_{1}^{(B^{2})}$ (54)
$\sigma_{xx}^{(B^{2},\Omega)}(\omega)$ $=$ $\frac{\mathit{v}_{F}^{3}}{120\pi\mu^{2}}\sigma_{1}^{(B^{2})}$ (55)
for J=2
$\sigma_{zz}^{(B^{2},\Omega)}(\omega)$ $=$ $\frac{5\mathit{v}_{F}\alpha_{2}^{2}}{32\mu\hbar}\sigma_{1}^{(B^{2})}$ (56)
$\sigma_{xx}^{(B^{2},\Omega)}(\omega)$ $=$ $\frac{\alpha_{2}^{2}\sigma_{1}^{(B^{2})}}{15\pi\mathit{v}_{F}\hbar^{2}}\Bigl{[}\frac{-2(15-25t_{s}^{2}+8t_{s}^{4})}{t_{s}^{6}}$ (57)
- $15\frac{(1-t_{s}^{2})^{2}}{t_{s}^{7}}\ln\frac{1-t_{s}}{1+t_{s}}\Bigr{]}$
for J=3
$\sigma_{zz}^{(B^{2},\Omega)}(\omega)=\frac{\mathit{v}_{F}\alpha_{3}^{\frac{2}{3}}\sigma_{1}^{(B^{2})}}{8\pi\mu^{\frac{2}{3}}\hbar^{\frac{4}{3}}}\Bigl{[}\frac{59049\Gamma{(\frac{11}{33})^{2}}{}_{2}F_{1}{(\frac{1}{3},\frac{11}{3},\frac{22}{3},\frac{2t_{s}}{-1+t_{s}})}}{34580{(\frac{1}{2}-\frac{t_{s}}{2})^{\frac{1}{3}}}\Gamma{(\frac{1}{3})}}\Bigr{]}$ (58)
$\begin{split}\sigma_{xx}^{(B^{2},\Omega)}(\omega)=&\frac{\mu^{\frac{2}{3}}\alpha^{\frac{4}{3}}\pi^{\frac{1}{2}}\sigma_{1}^{(B^{2})}}{8\mathit{v}_{F}\hbar^{\frac{8}{3}}}\frac{1}{{27{(3)^{\frac{1}{2}}}{(1-t_{s})^{\frac{2}{3}}}t_{s}^{4}{(-1+t_{s}^{2})}\Gamma{(\frac{5}{6})}\Gamma{(\frac{17}{3})}}}\bigg{[}6272(1+t_{s})\bigg{\{}\big{(}1190+t_{s}(1530+t_{s}(15-t_{s}(335+24t_{s}))\big{)}\\ &{}_{2}F_{1}{\Big{(}\frac{-1}{3},\frac{10}{3},\frac{14}{3},\frac{2t_{s}}{-1+t_{s}}\Big{)}}+5(1+t_{s})(-238+t_{s}(-34+t_{s}(53+3t_{s})))_{2}F_{1}{\Big{(}\frac{-1}{3},\frac{13}{3},\frac{14}{3},\frac{2t_{s}}{-1+t_{s}}\Big{)}}\bigg{\}}\bigg{]}\end{split}$ (59)
So, we can rewrite the above components follow the power laws $\sigma_{zz}^{(B^{2},\Omega)}(\omega)\propto\frac{\alpha_{J}^{\frac{2}{J}}}{\mu^{\frac{2}{J}}\hbar^{(2-\frac{2}{J})}}$.
@@
for J=1
$\sigma_{zz}^{(B^{2},m)}(\omega)$ $=$ $\frac{\mathit{v}_{F}^{3}}{120\pi\mu^{2}}\sigma_{1}^{(B^{2})}\Bigl{[}-3+5t_{s}^{2}\Bigr{]}$ (60)
$\sigma_{xx}^{(B^{2},m)}(\omega)$ $=$ $-\frac{\mathit{v}_{F}^{3}}{120\pi\mu^{2}}\sigma_{1}^{(B^{2})}$ (61)
for J=2
$\sigma_{zz}^{(B^{2},m)}(\omega)$ $=$ $\frac{\mathit{v}_{F}\alpha_{2}}{8\mu\hbar}\sigma_{1}^{(B^{2})}$ (62)
$\sigma_{xx}^{(B^{2},m)}(\omega)$ $=$ $\frac{2\alpha_{2}^{2}}{5\pi\hbar^{2}\mathit{v}_{F}}\sigma_{1}^{(B^{2})}\Bigl{[}\frac{(30-20t_{s}^{2}+t_{s}^{4})}{t_{s}^{6}}$ (63)
$+$ $\frac{5}{2}\frac{(6-6t_{s}^{2}+t_{s}^{4})}{t_{s}^{7}}\ln\frac{1-t_{s}}{1+t_{s}}]$
for J=3
@@
$\sigma_{zz}^{(B^{2},m)}(\omega)=\frac{\mathit{v}_{F}\alpha_{3}^{\frac{2}{3}}\sigma_{1}^{(B^{2})}}{8\pi\mu^{\frac{2}{3}}\hbar^{\frac{4}{3}}}\frac{2^{\frac{1}{3}}(-(-1+t_{s})^{\frac{1}{3}}\Gamma{(\frac{-1}{3})^{2}}}{{729t_{s}^{5}(-(-1+t_{s}))^{\frac{2}{3}}\Gamma{(\frac{7}{3})}}}\bigg{[}-\big{\{}728+3t_{s}\big{(}104+3t_{s}(-117+t_{s}(-41+3t_{s}(1+t_{s})(15+2t_{s})))\big{)}\big{\}}\\ {}_{2}F_{1}{\Big{(}\frac{1}{3},\frac{5}{3},\frac{7}{3},\frac{2t_{s}}{-1+t_{s}}\Big{)}}-{(1+t_{s})\big{\{}-728+3t_{s}\big{(}-208+t_{s}(143+3t_{s}(34+t_{s}(-11+6t_{s})))\big{)}\big{\}}}{}_{2}F_{1}{\Big{(}\frac{4}{3},\frac{5}{3},\frac{7}{3},\frac{2t_{s}}{-1+t_{s}}\Big{)}}\Bigr{]}$ (64)
@@
$\sigma_{xx}^{(B^{2},m)}(\omega)=\frac{\mu^{\frac{2}{3}}\alpha_{3}^{\frac{4}{3}}\sigma_{1}^{(B^{2})}}{8\pi^{\frac{1}{2}}\hbar^{\frac{8}{3}}\mathit{v}_{F}}\frac{-16\Gamma{(\frac{7}{6})}}{5(3)^{\frac{1}{2}}(1-t_{s})^{\frac{2}{3}}t_{s}^{6}(-1+t_{s}^{2})\Gamma{(\frac{5}{3})}}\bigg{[}\big{\{}5236+t_{s}\big{(}1309+t_{s}(-6762+t_{s}(-1393+2t_{s}(982+\\ 3(41-10t_{s})t_{s})))\big{)}\big{\}}{}_{2}F_{1}{\Big{(}\frac{-1}{3},\frac{4}{3},\frac{8}{3},\frac{2t_{s}}{-1+t_{s}}\Big{)}}+4(-1309+1631t_{s}^{2}-442t_{s}^{4}+12t_{s}^{6}){{}_{2}F_{1}{\Big{(}\frac{-1}{3},\frac{7}{3},\frac{8}{3},\frac{2t_{s}}{-1+t_{s}}\Big{)}}}\Big{]}$ (65)
where $\sigma_{1}^{(B^{2})}=\frac{e^{4}\tau B^{2}}{\pi\hbar(1-i\omega t)}$.
So, we can rewrite the above components follow the power laws $\sigma_{zz}^{(B^{2},m)}(\omega)\propto\frac{\alpha_{J}^{\frac{2}{J}}}{\mu^{\frac{2}{J}}}$ and $\sigma_{xx}^{(B^{2},m)}(\omega)\propto\frac{\alpha_{J}^{\frac{4}{J}}}{\mu^{\frac{2}{J}}\hbar^{(2-\frac{2}{J})}}$.
Obviously, from Eqs.(55) and (61), $\sigma_{xx}^{(B^{2},\Omega)}(\omega)$ and $\sigma_{xx}^{(B^{2},m)}(\omega)$, thus the total conductivity $\sigma_{xx}^{(B^{2})}(\omega)$ is equal to zero in the case of single-WSM [22]. For double WSM the total conductivity is enhanced(suppressed) below(above) the critical value $t_{0}$ of the tilt parameter as shown in Fig.(5a). In case of triple WSMs, the total conductivity is enhanced due to the orbital contribution(see Fig.(5b). Further, in the case of single WSM, the total conductivity component $\sigma_{zz}^{(B^{2})}(\omega)$ is suppressed(enhanced) below(above) the value of $t_{0}$ as shown in Fig.(6a)(see ref.[22]). For double WSM, the conductivity component $\sigma_{zz}^{(B^{2})}(\omega)$ the total conductivity is enhanced with tilt parameter $t_{0}$ as shown in Fig.(6b). For triple-WSM, the conductivity component $\sigma_{zz}^{(B^{2})}(\omega)$ the total conductivity is suppressed with tilt parameter $t_{0}$ as shown in Fig.(6c).
Refer to caption  Refer to caption
Figure 5: The dependence of the optical conductivity on the tilt $t_{+}=0.5$ for the case of B $\parallel t_{s}$. for (a) double WSM and (b) triple WSM. The other parameters are the same as those of Fig.(1)
Case-II In this case both longitudinal components as well as planer Hall conductivities are non-zero. The longitudinal components have the following expressions
for J=1
$\sigma_{xx}^{(B^{2},\Omega)}(\omega)$ $=$ $\frac{\mathit{v}_{F}^{3}\sigma_{1}^{(B^{2})}}{120\pi\mu^{2}}\Bigl{[}(8+13t_{s}^{2})\cos^{2}\gamma+\sin^{2}\gamma\Bigr{]}$ (66)
$\sigma_{zz}^{(B^{2},\Omega)}(\omega)$ $=$ $\frac{\mathit{v}_{F}^{3}}{120\pi\mu^{2}}\sigma_{1}^{(B^{2})}(1+7t_{s}^{2})$ (67)
for J=2
$\sigma_{xx}^{(B^{2},\Omega)}(\omega)$ $=$ $\frac{\alpha_{2}\mathit{v}_{F}}{256\pi^{2}\mu\hbar^{2}}\sigma_{1}^{(B^{2})}[(31+16t_{s}^{2})\cos^{2}\gamma+5\sin^{2}\gamma]$
$\sigma_{zz}^{(B^{2},\Omega)}(\omega)$ $=$ $\frac{\mathit{v}_{F}^{3}\sigma_{1}^{(B^{2})}}{60\pi\mu^{2}}(1+7t_{s}^{2})$ (69)
for J=3
$\begin{split}\sigma_{xx}^{(B^{2},\Omega)}(\omega)&=\frac{\mathit{v}_{F}\alpha_{3}^{\frac{2}{3}}\sigma_{1}^{(B^{2})}}{8\pi\hbar^{\frac{4}{3}}\mu^{\frac{2}{3}}}\Bigg{[}\cos^{2}\gamma\frac{1}{670208(1-t_{s})^{\frac{1}{3}}t_{s}^{4}}27\pi\bigg{\{}\big{(}4860-10917t_{s}^{2}+74328t_{s}^{4}+30464t_{s}\big{)}{}_{2}F_{1}{\Big{(}\frac{1}{3},\frac{2}{3},3,\frac{2t_{s}}{-1+t_{s}}\Big{)}}+(1+t_{s})\\ &\big{(}-40860+7677t_{s}^{2}+11288t_{s}^{4}\big{)}{{}_{2}F_{1}\big{(}\frac{4}{3},\frac{3}{2},3,\frac{2t_{s}}{-1+t_{s}}\big{)}}\bigg{\}}+\sin^{2}\gamma\frac{59049\Gamma(\frac{11}{3})^{2}{{}_{2}F_{1}\Big{(}\frac{1}{3},\frac{11}{3},\frac{22}{3},\frac{2t_{s}}{-1+t_{s}}\big{)}}}{276640\big{(}\frac{1}{2}-\frac{t_{s}}{2}\Big{)}^{\frac{1}{3}}\Gamma\big{(}\frac{1}{3}\big{)}}\bigg{]}\end{split}$ (70)
$\sigma_{zz}^{(B^{2},\Omega)}(\omega)=\frac{\mathit{v}_{F}^{3}\sigma_{1}^{(B^{2})}}{40\pi\mu^{2}}(1+7t_{s}^{2})$ (71)
So, we can rewrite the above components follow the power laws $\sigma_{zz}^{(B^{2},\Omega)}(\omega)\propto\frac{J}{\mu^{2}}$ and $\sigma_{xx}^{(B^{2},\Omega)}(\omega)\propto\frac{\alpha_{J}^{\frac{2}{J}}}{\mu^{\frac{2}{J}}\hbar^{2(1-\frac{1}{J})}}$.
@@
for J=1
$\sigma_{xy}^{(B^{2},\Omega)}(\omega)=\frac{\mathit{v}_{F}^{3}}{120\pi\mu^{2}}\sigma_{1}^{(B^{2})}(7+13t_{s}^{2})\sin\gamma\cos\gamma$ (72)
for J=2
$\sigma_{xy}^{(B^{2},\Omega)}(\omega)=\frac{\mathit{v}_{F}\alpha_{2}}{128\hbar\mu}\sigma_{1}^{(B^{2})}(13+8t_{s}^{2})\sin\gamma\cos\gamma$ (73)
for J=3
@@
$\sigma_{xy}^{(B^{2},\Omega)}(\omega)$ $=$ $\frac{\mathit{v}_{F}\alpha_{3}^{\frac{2}{3}}\sigma_{1}^{(B^{2})}}{8\pi^{\frac{3}{2}}\mu^{\frac{2}{3}}\hbar^{\frac{4}{3}}}\sin\gamma\cos\gamma\Biggl{[}\frac{1}{189(1-t_{s})^{\frac{1}{3}}}{\bigg{\{}15309(-1+t_{s})^{2}\Gamma{\Big{(}\frac{2}{3}\Big{)}}\Gamma{\Big{(}\frac{5}{6}\Big{)}}{{}_{2}F_{1}{\Big{(}\frac{1}{3},\frac{5}{3},\frac{7}{3},\frac{2t_{s}}{-1+t_{s}}\Big{)}}}}$ (74)
$+$ $\frac{1}{t_{s}^{4}\Gamma{(\frac{1}{6})\Gamma{(\frac{1}{3})}}}2{(3)^{\frac{1}{2}}}\pi^{2}\big{\{}\big{(}-130-3t_{s}(52+9t_{s}(-6+t_{s}(-8+t_{s}(287+9t_{s}(-74+35t_{s})))))\big{)}{}_{2}F_{1}{\Big{(}\frac{1}{3},\frac{8}{3},\frac{10}{3},\frac{2t_{s}}{-1+t_{s}}\Big{)}}$
$+$ $(1+t_{s})\big{(}130+9(-1+t_{s})t_{s}(-26+3t_{s}(-10-6t_{s}+45t_{s}^{2}))\big{)}{}_{2}F_{1}{\Big{(}\frac{4}{3},\frac{8}{3},\frac{10}{3},\frac{2t_{s}}{-1+t_{s}}\Big{)}}\big{\}}\bigg{\}}\Biggr{]}$
So, we can rewrite the above components follow the power laws $\sigma_{xy}^{(B^{2},\Omega)}(\omega)\propto\frac{\alpha_{J}^{\frac{2}{J}}}{\mu^{\frac{2}{J}}\hbar^{2(1-\frac{1}{J})}}$.
Refer to caption  Refer to caption  Refer to caption
Figure 6: The dependence of the optical conductivity on the tilt $t_{+}=0.5$ for the case of B $\perp t_{s}$. for (a) single WSM (b)double WSM and (c) triple WSM. The other parameters are the same as those of Fig.(1)
Refer to caption  Refer to caption  Refer to caption
Figure 7: The dependence of the optical conductivity on the tilt $t_{+}=0.5$ for the case of B $\perp t_{s}$. for (a) single WSM (b)double WSM and (c) triple WSM. The other parameters are the same as those of Fig.(1)
Refer to caption  Refer to caption  Refer to caption
Figure 8: The dependence of planar Hall conductivity on the tilt $t_{+}=0.5$ and $\gamma=\frac{\pi}{4}$ for (a) single WSM (b)double WSM and (c) triple WSM. The other parameters are the same as those of Fig.(1)
The magnetic orbital moment contributions of conductivities
for J=1
$\sigma_{xx}^{(B^{2},m)}(\omega)$ $=$ $\frac{\mathit{v}_{F}^{3}}{120\pi\mu^{2}}\sigma_{1}^{(B^{2})}\Bigl{[}(-3-6t_{s}^{2})\cos^{2}\gamma-\sin^{2}\gamma\Bigr{]}$
$\sigma_{zz}^{(B^{2},m)}(\omega)$ $=$ $\frac{\mathit{v}_{F}^{3}}{120\pi\mu^{2}}\sigma_{1}^{(B^{2})}(-1+t_{s}^{2})$ (76)
$\sigma_{xy}^{(B^{2},m)}(\omega)$ $=$ $\frac{\mathit{v}_{F}^{3}}{120\pi\mu^{2}}\sigma_{1}^{(B^{2})}(-2-6t_{s}^{2})\sin\gamma\cos\gamma$ (77)
for J=2
$\sigma_{xx}^{(B^{2},m)}(\omega)$ $=$ $\frac{\mathit{v}_{F}\alpha_{2}}{64\mu\hbar}\sigma_{1}^{(B^{2})}(2\cos^{2}\gamma-\sin^{2}\gamma)$ (78)
$\sigma_{zz}^{(B^{2},m)}(\omega)$ $=$ $\frac{\mathit{v}_{F}^{3}}{60\pi\mu^{2}}\sigma_{1}^{(B^{2})}(-1+t_{s}^{2})$ (79)
$\sigma_{xy}^{(B^{2},m)}(\omega)$ $=$ $\frac{\mathit{v}_{F}\alpha_{2}}{32\hbar\mu}\sin\gamma\cos\gamma\sigma_{1}^{(B^{2})}$ (80)
for J=3
@@
$\begin{split}\sigma_{xx}^{(B^{2},m)}(\omega)&=\frac{\mathit{v_{F}}\alpha_{3}^{\frac{2}{3}}}{8\pi^{\frac{1}{2}}\hbar^{\frac{4}{3}}\mu^{\frac{2}{3}}}\sigma_{1}^{(B^{2})}\Biggl{[}-\frac{1}{24(1-t_{s})^{\frac{1}{3}}t_{s}^{5}\Gamma{(\frac{1}{6})}}\Gamma{(\frac{5}{3})}\bigg{\{}(1-t_{s})_{2}F_{1}{\Bigl{(}\frac{4}{3},\frac{5}{3},\frac{7}{3},\frac{2t_{s}}{-1+t_{s}}\Bigr{)}}\Big{(}6(-91+2t_{s}(-39+t_{s}(-13+\\ &3t_{s}(-5+2t_{s}(5+3t_{s})))))\cos^{2}\gamma+\big{(}-182+3t_{s}(-52-9(-3+t_{s})t_{s}(1+t_{s})))\big{)}\sin^{2}\gamma\Big{)}+_{2}F_{1}{\Bigl{(}\frac{1}{3},\frac{5}{3},\frac{7}{3},\frac{2t_{s}}{-1+t_{s}}\Bigr{)}}\\ &\big{\{}6(91+t_{s}(39+2t_{s}(-26+3t_{s}(-2+3t_{s}(-5+2(-2+t_{s})t_{s}))))\big{\}}\cos^{2}\gamma+(182+3t_{s}(26+t_{s}(-79\\ &+27t_{s}(-1+t_{s}+t_{s}^{2}))))\sin^{2}\gamma\big{\}}\bigg{\}}\Bigg{]}\end{split}$ (81)
$\sigma_{zz}^{(B^{2},m)}(\omega)=\frac{\mathit{v}_{F}^{3}\sigma_{1}^{(B^{2})}}{40\pi\mu^{2}}(-1+t_{s}^{2})\\$ (82)
$\begin{split}\sigma_{xy}^{(B^{2},m)}(\omega)&=\frac{\mathit{v}_{F}\alpha_{3}^{\frac{2}{3}}\sigma_{1}^{(B^{2})}}{8\pi^{\frac{3}{2}}\mu^{\frac{2}{3}}\hbar^{\frac{4}{3}}}{\sin\gamma\cos\gamma}\frac{\Gamma\big{(}\frac{-1}{3}\big{)}\Gamma\big{(}\frac{-1}{6}\big{)}}{1296(1-t_{s})^{\frac{1}{3}}t_{s}^{5}}\bigg{[}\big{\{}-364+3t_{s}\big{(}-52+t_{s}(25-3t_{s}(1+3t_{s}-(23+\\ &t_{s}(-19+8t_{s}))))\big{)}\big{\}}{{}_{2}F_{1}\big{(}\frac{1}{3},\frac{5}{3},\frac{7}{3},\frac{2t_{s}}{-1+t_{s}}\big{)}}-(1+t_{s})\big{\{}-384+3t_{s}\big{(}-104+t_{s}(-79+3t_{s}(-26+\\ &t_{s}(43+24t_{s})))\big{)}\big{\}}{{}_{2}F_{1}\big{(}\frac{4}{3},\frac{5}{3},\frac{7}{3},\frac{2t_{s}}{-1+t_{s}}\big{)}}\bigg{]}\end{split}$ (83)
So, we can rewrite the above components follow the power laws $\sigma_{zz}^{(B^{2},m)}(\omega)\propto\frac{J}{\mu^{2}}$ , $\sigma_{xx}^{(B^{2},m)}(\omega)\propto\frac{\alpha_{J}^{\frac{2}{J}}}{\mu^{\frac{2}{J}}\hbar^{2(1-\frac{1}{J})}}$ and $\sigma_{xy}^{(B^{2},m)}(\omega)\propto\frac{\alpha_{J}^{\frac{2}{J}}}{\mu^{\frac{2}{J}}\hbar^{2(1-\frac{1}{J})}}$.
Refer to caption  Refer to caption  Refer to caption
Figure 9: The frequency dependence of optical conductivity at tilt $t_{+}=0.5$, $\gamma=\frac{\pi}{4}$ and B = 1 T. for (a) single WSM (b)double WSM and (c) triple WSM.The other parameters are the same as those of Fig.(1)